The humane society reports that of 428 animals at their local animal shelter, 376 are household pets and the remaining 52 are wildlife animals. Over the weekend, 29 of the household pets were adopted and 10 of the wild life animals were released back into the wild. Is there sufficient evidence to indicate a difference in the number of animals leaving the shelter over the weekend for the two types (household pets and wildlife animals)? use the p-value approach at the 1% level of significance.
Answer #1

H0 : p1 = p2
Ha: p1 not equals to p2

test statistics:

pcap = 29 +10/428
= 0.0911

p1 = 29/376 =0.0771

p2 = 10/52 = 0.1923
z = (p1 -p2)/sqrt(pcap*(1-pcap)*(1/n1+1/n2)
= (0.0771 - 0.1923)/sqrt(0.0911 *(1-0.09110/*(1/376 + 1/52))
= -2.7051

p value = .0068

Reject H0